The Gulf Power Company currently is an all-equity firm. The value of Gulf Power's equity is $12,000,000 and there are 600,000 shares outstanding. The expected annual EBIT of Gulf Power is $2,400,000. Those earnings are also expected to remain constant into the foreseeable future. Gulf Power is in the 40-percent tax bracket. The Gulf Power Company plans to announce that it will issue $3,000,000 of perpetual bonds and uses the proceeds to repurchase common stock. The bonds will have a 5-percent coupon rate. After the sale of the bonds, Gulf Power will maintain the new capital structure indefinitely. The MM theory applies.   

What is the firm's cost of capital before the capital restructuring?

What is the number of shares repurchased?

What is the firm's cost of capital after the capital restructuring?

What is Gulf Power's cost of equity after the capital restructuring?

What is the firm's net income (NI) after the capital restructuring?
Answer #1

Ans:- Expected EBIT of Gulf Power is $2400000. Tax is given 40%. Value of Gulf Power equity is 12000000 and there are 600000 outstanding shares.

Price Per share of Gulf Power will be (Value of Gulf Power equity/ outstanding shares) = 12000000/600000 = 20.

(a) Note:- Firms cost of capital is the rate of return that the company would have earned by putting their money in different investments with the same risk. Capital restructuring is nothing but when the firm changes the mixture of debt and equity its capital structure.

Now first we need to find net income after tax.

Net Income after tax for Gulf Power will be given by EBIT*(1-Tax rate) =2400000* (1-0.40) = 1440000.

Return on equity for Gulf Power will be given by( Income after-tax/ Value of equity)

(1440000/12000000)*100 =12%

(b) To find the number of shares repurchased first we need to find the value of the firm after the announcement of Perpetual Bonds or debts.

Value of Gulf power will be given by ( Value of Gulf power equity + Debt* Tax rate )

= 12000000 + 3000000*40% = 13200000.

  Now the price per share after the announcement will be given by(Value of firm after announcement/ outstanding shares)

= 13200000/600000 = 22.

Now the number of shares repurchased will be (Value of debt issued by firm/ Price per share)

= 3000000/22 = 136364.(approx).

(d) In this part, we need to find the cost of equity capital after the restructuring

Value of Debt of Gulf Power is 3000000.

Value of Equity of Gulf power will be 13200000 - 3000000 = 10200000

Now the cost of equity capital after the recapitalization will be given by

= rate of equity return + (debt/equity) * ( rate of equity return - Bond equity rate)*(1- Tax rate)

= 12% + (3000000/102000000)* (12% - 5%)* (1-0.4) =14.01%

(c) Now the cost of capital after capital restructuring. In this part we need to find the weighted average cost of capital i.e (WACC)  

Weighted Average Cost of Capital will be given by

(Value of equity/ Total market value of firm)* Cost of equity+ (Value of Debt/ Total Market Value of firm)*cost of debt *(1-Tax)

= (10200000/13200000)*14.01% + (3000000/13200000) * 5%* (1-0.40) = 11.50%.

(e) In this part, we need to find the Net Income after the capital restructuring.

Earning Before Interest and Tax is 2400000 - Interest i.e 2400000 - 5% of 3000000 =2250000

EBT = 2250000 *(1 - Tax) = 2250000*(1-0.40)= 1350000 is the required Net Income.